Q5b. If A=5+2i and B=8+6i, find C = (A-B)+(A) B i) D= ((A/B) +B) + (A – B)
Answer #1

Co (5+272(8759) + (5+21) 8+6i = (28tuci) + (5+2;) - stoi = 33 + 481 8tci Rationalize = 33+481 1 8-611 8+61 -64 turi 64 + 36 =ii) Dec (1+B)* CA-6] le * 8+i) (542-8--7) *** ***2*F} ](3-4;) =rstzit (8+61)(8+61) 7 i 8f6i ui) = (strit 28+9677 Le &to 41) (- - %۱۰ 7 )5 - ) - کاک - 212 - + و ) ( 8 - 17) - Cb و )( 8- 49) = ( ح ط